OpenStudy (anonymous):

If sinA = 2/3 find cos A/2 =? need help ASAP

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

first find \(\cos(A)\)

OpenStudy (misty1212):

then use \[\cos(\frac{A}{2})=\sqrt{\frac{1+\cos(A)}{2}}\]

OpenStudy (anonymous):

HOW to get cos A?

OpenStudy (misty1212):

draw a triangle

OpenStudy (misty1212):



OpenStudy (misty1212):

there is a picture on an angle whose sine is \(\frac{2}{3}\) find the third side via pythagoras

OpenStudy (misty1212):

once you have the missing side, you know the cosine let me know when you get it

OpenStudy (anonymous):

Oh i get it ... i'll give it a try... with some other exercises

OpenStudy (anonymous):

Waiting for lynfran's reply